What part of the general slope-intercept equation represents the slope?
What part of the general slope-intercept equation represents the y-intercept?
Define the y-intercept.
Define the x-intercept.
Identify the x-intercepts and y-intercepts of the equations for problems #6 and #7. Describe how to graph the line.
#6. 2x + y = 13
#7. -16x + y = -4
Describe how you found the x-intercepts and y-intercepts in problems #6 and #7.

Answer by checkley75(3666)

2X+Y=13, Y=-3X+13 SETTING X=0 & SOLVE FOR Y, Y=13 OR (0,13) SETTING Y=0 & SOLVE FOR X, 0=-3X+13, -3X=-13, X=-13/-3, X=13/3 OR (13/3,0)
-----------------------------------------------------------------
-16X+Y=-4, Y=16X-4 SETTING X=0 & SOLVE FOR Y, Y=-4 OR (0,-4) SETTING Y=0 & SOLVE FOR X, 0=16X-4, 16X=4, X=4/16, X=1/4 OR (1/4,0)